# Proposal for a 3-Wave Mixing Element with Quantum Paraelectric Materials

### v1

@@ Appendix D Nanofabrication
Figure 6: Example PANDA designs. (a) A parallel plate capacitor is made from a thin layer STO or KTO sandwiched by bottom and top electrode layers. This structure sits on top of a conventional substrate, for example silicon or sapphire. Geometric inductance from the wiring layer combines with the parallel plate capacitance to form a resonant circuit. The resonator is connected to a port using an inductive coupling element, e.g. a stepped-impedance filter as in Refs. [21, 30, 24], which sets the external coupling rate $\kappa_{\mathrm{ext}}$. (b) Alternatively, a PANDA may be formed using a coplanar waveguide (CPW) resonator, like the device in Ref. [40], where a CPW of impedance $Z_{\mathrm{CPW}}$ is connected to a $Z_{p}=50\,\mathrm{\Omega}$ port. Because the CPW is fabricated on STO or KTO substrate, permittivity is high and $Z_{\mathrm{CPW}}$ can be very low compared to $50\penalty 10000\ \,\mathrm{\Omega}$. This creates an impedance mismatch which sets the external coupling rate $\kappa_{\mathrm{ext}}$, forming a half-wavelength resonator. In contrast to a conventional CPW designed to have impedance of $50\,\mathrm{\Omega}$ , the gap (see inset) should be as small as possible compared to the width, to further reduce the impedance and to minimize the bias voltage needed to tune permittivity. In both (a) and (b), superconducting material (e.g. Nb) can be used to minimize loss and the conductor should be as thick as possible to minimize kinetic inductance. Charge on the center conductor can be controlled via a bias-tee.
@@
Our proposed design is a lumped element resonator formed by a parallel plate capacitor shunted by an external inductance, as in Ref. [55]. A co-planar waveguide (CPW) structure [40], shown for example om Fig. 6, is able to leverage the same physics but with a weaker field confinement. In addition, the spatial dependence of the electric field within the CPW substrate complicates our theoretical treatment, because the relative orientation of the DC and RF electric fields will play a role in how much of the nonlinearity is activated. This effect is minimized for our lumped element capacitor design.
@@
To create the requisite parallel capacitor, thin films of STO or KTO can be directly grown on a variety of substrates, even including directly on a metallic superconducting layer [56, 57]. Likely, homoepitaxial growth will not be a viable route due to the large dielectric constant of the substrate. Depending on the substrate, epitaxial mismatch strain (i.e. with silicon substrates) will degrade the low temperature dielectric properties. This can result in a lower tunability and alter the optimal operating temperature of the device.
@@
Oxide films can also be grown on water soluble buffer layers, allowing for layer transfer of the epitaxial thin films [58]. Remote epitaxy on materials like graphene also can produce thin films [59]. Heterogeneous bonding and removal of the growth substrate (i.e. Silicon) is an alternate route to transfer films. Finally, bulk thinning of crystals is another potential pathway to create low-strain, bulk-like oxide thin films of a variety of substrates. Bonding methods can be interface-free and room temperature, compatible with existing devices on the substrate. Finally, to optimize performance, the tunability of the dielectric constant with field can be optimized with precise strain, doping, and isotope control. The exact coefficient of thermal expansion mismatch with the substrate will be critical.
@@
For microwave devices, the oxide thin film needs to be laterally small and thin. The film would be transferred onto a conventional, low-loss substrates (e.g. sapphire or silicon) with a superconducting layer deposited on top. The resulting film on the superconductor-substrate stack would then be patterned into the desired capacitor size through dry or wet etching. The top surface would then be metalized with another superconducting layer. The bottom superconducting layer would connect to the circuit ground, while the top layer would be connected by a superconducting crossover to the substrate that leads to the on-chip meander inductor. As a result, small patterned coupons of thin film oxides can be placed where they are desired on a standard superconducting device substrate and integrated seamlessly with existing circuitry. Flip-chip bonding techniques can also be used to place capacitors as desired on a microwave circuit [41].
@@
While oxide thin films can be grown and transferred at the wafer scale, little is known about their low temperature dielectric performance at high frequencies, requiring further study. Likely, thin film properties will be degraded from the bulk. There may also be significant “dead layer” effects which limit the applicability of very thin films. Despite this, due to proximity to the quantum phase transition, the dielectric constant will certainly still remain very large and tunable. Even with degraded performance from realistic materials, such devices may still constitute a competitive alternative to other cryogenic parametric mixing elements.

### v2

@@ Appendix E Nanofabrication
Figure 7: Example PANDA designs. (a) A parallel-plate capacitor is made from a thin layer STO or KTO sandwiched by bottom and top electrode layers. This structure sits on top of a conventional substrate, for example silicon or sapphire. Geometric inductance from the wiring layer combines with the parallel-plate capacitance to form a resonant circuit. The resonator is connected to a port using an inductive coupling element, e.g. a stepped-impedance filter as in Refs. [21, 30, 24], which sets the external coupling rate $\kappa_{\mathrm{ext}}$. (b) Alternatively, a PANDA may be formed using a coplanar waveguide (CPW) resonator, like the device in Ref. [44], where a CPW of impedance $Z_{\mathrm{CPW}}$ is connected to a $Z_{p}=50\,\mathrm{\Omega}$ port. Because the CPW is fabricated on STO or KTO substrate, permittivity is high and $Z_{\mathrm{CPW}}$ can be very low compared to $50\penalty 10000\ \,\mathrm{\Omega}$. This creates an impedance mismatch which sets the external coupling rate $\kappa_{\mathrm{ext}}$, forming a half-wavelength resonator. In contrast to a conventional CPW designed to have impedance of $50\,\mathrm{\Omega}$ , the gap (see inset) should be as small as possible compared to the width, to further reduce the impedance and to minimize the bias voltage needed to tune permittivity. In both (a) and (b), superconducting material (e.g. Nb) can be used to minimize loss and the conductor should be as thick as possible to minimize kinetic inductance. Charge on the center conductor can be controlled via a bias tee.
@@
Our proposed design is a lumped-element resonator formed by a parallel plate capacitor shunted by an external inductance, as in Ref. [61]. A co-planar waveguide (CPW) structure [44], shown for example in Fig. 7, is able to leverage the same physics but with a weaker field confinement. In addition, the spatial dependence of the electric field within the CPW substrate complicates our theoretical treatment, because the relative orientation of the dc and rf electric fields will play a role in how much of the nonlinearity is activated. This effect is minimized for our lumped-element capacitor design.
@@
To create the requisite parallel capacitor, thin films of STO or KTO can be directly grown on a variety of substrates, even including directly on a metallic superconducting layer [62, 63]. Likely, homoepitaxial growth will not be a viable route due to the large dielectric constant of the substrate. Depending on the substrate, epitaxial mismatch strain (i.e. with silicon substrates) will degrade the low-temperature dielectric properties. This can result in a lower tunability and alter the optimal operating temperature of the device.
@@
Oxide films can also be grown on water-soluble buffer layers, allowing for layer transfer of the epitaxial thin films [64]. Remote epitaxy on materials like graphene also can produce thin films [65]. Heterogeneous bonding and removal of the growth substrate (i.e. silicon) is an alternate route to transfer films. Finally, bulk thinning of crystals is another potential pathway to create low-strain, bulklike oxide thin films of a variety of substrates. Bonding methods can be interface-free and room temperature, compatible with existing devices on the substrate. Finally, to optimize performance, the tunability of the dielectric constant with field can be optimized with precise strain, doping, and isotope control. The exact coefficient of thermal expansion mismatch with the substrate will be critical.
@@
For microwave devices, the oxide thin film needs to be laterally small and thin. The film would be transferred onto a conventional, low-loss substrates (e.g. sapphire or silicon) with a superconducting layer deposited on top. The resulting film on the superconductor-substrate stack would then be patterned into the desired capacitor size through dry or wet etching. The top surface would then be metalized with another superconducting layer. The bottom superconducting layer would connect to the circuit ground, while the top layer would be connected by a superconducting crossover to the substrate that leads to the on-chip meander inductor. As a result, small patterned coupons of thin-film oxides can be placed where they are desired on a standard superconducting device substrate and integrated seamlessly with existing circuitry. Flip-chip bonding techniques can also be used to place capacitors as desired on a microwave circuit [45].
@@
While oxide thin films can be grown and transferred at the wafer scale, little is known about their low-temperature dielectric performance at high frequencies, requiring further study. Likely, thin-film properties will be degraded from the bulk. There may also be significant “dead layer” effects, which limit the applicability of very thin films. Despite this, due to the proximity to the quantum phase transition, the dielectric constant will certainly still remain very large and tunable. Even with degraded performance from realistic materials, such devices may still constitute a competitive alternative to other cryogenic parametric-mixing elements.